# Speak the Same Language: Global LiDAR Registration on BIM Using Pose Hough Transform

### v1

@@ IV-C Triangle Descriptors
Upon preprocessing the BIM and the LiDAR submap, we employ triangle descriptors to establish correspondences between them. This method is inspired by the STD introduced in [28]. We leverage both corners and wall segments to construct these descriptors.
Corners extracted from the LiDAR submap and subBIM are used to form a neighborhood graph, where vertices represent corners and edges connect two corners if they are within a maximum distance of $L_{max}$. Cliques in this graph with three vertices are used to form triangle descriptors, denoted by $\Delta$. These triangles are formed by corner triplets $(A,B,C)$, as illustrated in Figure 4. The attributes of a triangle descriptor $\Delta$ include:
@@
These sorted side lengths and angles capture both the geometric and structural details of the local environment, distinguishing this approach from the original STD descriptor [28]. As shown in Figure 5, querying each triangle in the submap can yield multiple responses in the subBIM due to repetitive architectural patterns. The inclusion of angles $\alpha,\beta,\gamma$ helps to differentiate triangles that have identical shapes but differ in their angular relationships with adjacent wall segments, thus significantly reducing incorrect correspondences.
For efficient data management and retrieval, we construct a hash table for both LiDAR submap and subBIMs. The keys are formed using the triangle descriptor $\Delta=[\|AB\|,\|BC\|,\|AC\|,\alpha,\beta,\gamma]$, with $\Delta\in\mathbf{R}^{6}$, and are quantized based on side length resolution $r_{\mathrm{s}}$ and angle resolution $r_{\mathrm{a}}$ as detailed in Table II and analyzed in Section VI-C1. Multiple corner triplets that yield the same hash key are stored together:
$\Delta\mathcal{S}$ =$=\{\Delta_{i}^{s}=\{(A,B,C)_{j}\}_{j=1}^{K_{i}^{s}}\}_{i=1}^{N^{s}},$
$\Delta\mathcal{B}$ =$=\{\Delta_{i}^{b}=\{(A,B,C)_{j}\}_{j=1}^{K_{i}^{b}}\}_{i=1}^{N^{b}},$
where $N^{s}$ and $N^{b}$ denote the number of triangle descriptors, and $K^{s}$ and $K^{b}$ represent the number of corner triplets for each descriptor in the submap and subBIM, respectively.
During the subBIM retrieval stage, hash keys from the submap are used to vote for potential matched subBIMs. The retrieval score for a subBIM is determined by the number of key intersections with the submap’s keys. This process is expedited by the $O(1)$ query complexity of the hash table. Then we retain all subBIMs with a nonzero number of votes as potential matching candidates, given the presence of numerous repetitive patterns that generate significant ambiguity in indoor scenes. For each subBIM candidate, the matched triangle descriptor can naturally be transformed into a set of corner triplet correspondences by exhaustively matching all corner triplets within two descriptors. Though triplet-based correspondences are with a extremely large number and high outlier ratio, they are crucial for the transformation estimation process, which will be discussed in the next section.
Figure 5: The figure demonstrates the effectiveness of triangle descriptors in matching results, with and without angle inclusion. On the left, a highlighted corner triplet from the LiDAR submap is shown in red, which has multiple matches from the subBIM using descriptor-based matching. Incorporating angle information significantly enhances the discriminative power of the triangle descriptor by encoding the local geometric structure around the triangle, which markedly reduces the false positives.
@@ V Back-end: Transformation Voting and Verification
This section outlines the methodology for deriving pose transformation candidates from subBIM candidates, utilizing a combination of the Hough transform and a voting mechanism. This process is particularly critical given the high ratio of outliers among corner triplet correspondences (Section V-A). Following this, we detail a verification scheme designed to identify the optimal transformation for aligning each submap with its corresponding BIM region, ensuring the reliable registration (Section V-B).
V-A Pose Hough Transformation and Voting
@@
The Hough transform is a robust technique in computer vision for converting raw measurements into a parameter space, where optimal parameters are identified through a voting mechanism to effectively reject outliers. Our implementation includes two primary steps:
Parameter Space Construction: The parameter space, based on the $\operatorname{SE}(2)$ Lie group, includes 2 degrees of freedom (DOF) for translation and one for yaw. This space is rasterized into voxels at resolutions specified by $r_{\mathrm{xy}}$ for x-y translation and $r_{\mathrm{yaw}}$ for yaw, as detailed in Table II and discussed in the ablation study (Section VI-C2).
Voting by Corner Triplet Correspondences: For each corner triplet correspondence $\left[(A,B,C)_{s},(A,B,C)_{b}\right]$, transformation parameters $(x,y,yaw)$ are computed using a closed-form solution for 2D registration [29]. A line segment overlap-based filter is applied to ensure that line segments intersecting at a submap corner overlap with those at the corresponding BIM corner. Only triplet correspondences that pass this filter contribute to the voting process.
For each subBIM, multiple corner triplet correspondences $\{\left(\Delta_{i}^{s},\Delta_{i}^{b}\right)\}_{i=1}^{N_{c}}$ are obtained through triangle descriptor based matching. Each corner triplet from $\Delta^{s}$ is matched against every triplet in $\Delta^{b}$, forming a minimum vote set where each voxel in the parameter space receives at most one vote. All minimum vote sets from $N_{c}$ descriptor correspondences are then combined to form the vote set for one subBIM candidate. In cases where voxels receive votes from multiple subBIM candidates, only the voxel with the maximum votes is retained. The final step involves grouping all pose voxels by their vote counts and selecting the top $J$ groups with the highest votes. The average votes falling into the same voxel represent our transformation candidate.
The computational complexity of this stage is primarily focused on the Hough transform step, denoted as $O(n)$, where $n$ is the number of corner triplet correspondences. This complexity is comparatively efficient, especially when juxtaposed against alternative transformation estimation methods such as RANSAC and TEASER++ (Section VI-C3).
@@ V-B Transformation Verification
Given transformation candidates for $J$ groups, our objective is to develop a verification scheme to determine the optimal transformation. While corners serve as a common reference between the submap and subBIMs in previous sections, they do not fully capture the geometric structure, leading to potential inaccuracies in alignment determination. To address this, we leverage walls, which encompass the majority of the common geometric structure between the submap and subBIMs, to verify the transformation candidates. The verification process consists of two components: wall preprocessing and similarity score computation.
In the wall preprocessing step, we project the wall point cloud data from both the subBIM and the transformed submap into images $I_{\mathcal{B}}$ and $I_{\mathcal{S}}$ at a pixel scale $s_{\mathrm{I}}$. However, directly comparing these images might not accurately reflect the true alignment quality for two primary reasons. First, the actual construction conditions (as-built data) may differ from the BIM designs (as-designed data). To mitigate this discrepancy, we apply a dilation to the wall pixels in the submap image by a radius of 1 meter. Second, as the submap only partially overlaps with the subBIM, we define a region of interest (ROI) within the subBIM image to focus the assessment on corresponding areas where the submap image exhibits wall pixels.
The similarity between the two images is quantified using the normalized cross correlation coefficient (NCC), a standard measure for image similarity:
$\mathtt{score}(I_{S},I_{B})=\frac{\sum_{u,v}\left(I_{S}(u,v)\cdot I_{B}(u,v)\right)}{\sqrt{\sum_{u,v}I_{S}(u,v)^{2}}\cdot\sqrt{\sum_{u,v}I_{B}(u,v)^{2}}}$ (3)
where $(u,v)$ are the coordinates within the ROI.
A typical case study illustrating this verification process is presented in Figure 6. Despite generating multiple transformation candidates through a Hough transform and voting procedure, the wall-based transformation verification effectively resolves ambiguities from the corner correspondences by selecting the transformation that best aligns the submap with the subBIM.
(a) More votes with lower similarity score
(b) Less votes with higher similarity score
Figure 6: This figure demonstrates the voting and verification process for a submap and subBIM pair, where (a) and (b) show the number of votes and similarity scores obtained by different transform candidates, respectively. The designed verification scheme could select the true transformation from multiple candidates based on the designed similarity score.
@@ VI Experiments
This section is structured as follows: In Section VI-A, we introduce the dataset, evaluation metrics, and implementation details of our proposed method, as well as those of the methods against which we compare. Section VI-B presents comparative experiments that assess the registration recall of our method relative to other advanced methods. In Section VI-C, we explore how different parameters affect the performance of our proposed algorithm through an ablation study. Section VI-D provides visualizations of our algorithm applied in real-world scenarios and discusses cases where the algorithm failed.
@@ VI-A1 LiBIM-UST Dataset
TABLE I: FusionBIM Dataset Information
Sensor Type Sequence Total Distance (m) Submap Distance (m) Number of Submaps
Ouster-128 Building Day 402.14 10 36
Mid-360 2f-office 533.78 15 40
30 15
We developed the LiBIM-UST dataset using the publicly available FusionPortable dataset [30] and additional data collected by our team for evaluation purposes. This dataset features the BIM of the main academic building at the Hong Kong University of Science and Technology (HKUST). It includes 91 LiDAR submaps and the ground truth pose transformations between them, as detailed in Table I.
The ”Building Day” sequence, part of the FusionPortable dataset [30], was captured in the lobby of the academic building using a mechanical rotating LiDAR OS1-128. The other was collected in office corridors using a solid-state LiDAR MID-360, which were gathered by our team. The sensor configuration for the self-collected data is depicted in Figure 7.
As described in Section IV-A1, LiDAR scans are accumulated over a period based on the submap distance $d_{s}$ to generate submaps. Given that different sequences have varying fields of view, we set distinct $d_{s}$ values for each, as indicated in Table I. To ensure adequate differentiation between submaps, their overlap ratio is limited to approximately 50%.
To determine the ground truth of the pose transformations from the submaps to the BIM, we initially align them manually and subsequently refine the alignments using the ICP algorithm.
TABLE II: Parameters of Our Method
Parameters Value
Line Detection Pixel scale $s_{\mathrm{I}}$ 12 pixels / m
Minmum length $L_{\mathrm{min}}$ 16 pixels
Triangle Descriptor Side length resolution $r_{\mathrm{s}}$ 1.25 m
Angle resolution $r_{\mathrm{a}}$ 3.8
Maximum side length $L_{\mathrm{max}}$ 37.5 m
Hough Transform x-y resolution $r_{\mathrm{xy}}$ 2.5 m
Yaw resolution $r_{\mathrm{yaw}}$ 5.0
TABLE III: Global Registration Evaluation on LiBIM-UST Dataset
Front-end Back-end Registration Recall(%) Average Time(s)
Building Day 2f-office Building Day 2f-office
$d_{s}$=10m $d_{s}$=15m $d_{s}$=30m $d_{s}$=10m $d_{s}$=15m $d_{s}$=30m
Image Matching- based Method SIFT [31] RANSAC 0.00 0.00 0.00 3.31 2.95 3.30
Superpoint [32]+Lightglue [33] RANSAC 0.00 0.00 0.00 2.24 2.31 2.33
Point Cloud- based Method FPFH [34] TEASER++ [35] 0.00 0.00 0.00 20.89 9.71 17.07
Voxel Map 3D BBS [36] 58.33 52.50 86.67 105.73 748.63 1122.47
Ours Triangle Descriptor Hough Voting 91.67 45.00 93.33 25.44 36.38 158.16
@@ VI-A3 Comparisons and Implementation
Image Registration: For image registration, intermediate images from the line extraction process (Section IV-B) are used. We implemented the SIFT [31] and a combination of Superpoint [32] with LightGlue [33], representing traditional and deep learning-based methods, respectively. RANSAC, with up to 4,000,000 iterations, was used for pose estimation.
Point Cloud Registration: We employed two common global point cloud registration methods: FPFH [34] combined with TEASER++ [35] and 3D-BBS [36], a branch-and-bound based method. For FPFH, we set the normal search radius to 1 m and the feature search radius to 2.5 m. For TEASER++, the noise bound is set at 0.4 m. For 3D-BBS, the parameter $\mathrm{min\_level\_res}$ is set to 1.0 m for the lobby and 0.5 m for office corridors. Additionally, recent works STD [28] and Outram [23], which share a similar correspondence generation module with our approach but differ in transformation solvers, are also compared. Their key point extraction and descriptor construction methods were not directly applicable; hence, we integrated only their backend into our framework and evaluated it in the ablation experiments (Section VI-C3).
Our Method: Detailed parameters of our method are listed in Table II. The decomposition of BIM and the Hash table configuration for each subBIM were predetermined. For simplicity and to expedite the verification process, our prototype was implemented in Python without multi-threading, affecting runtime efficiency. The experiments were conducted on an Intel i7-10700 CPU and an Nvidia RTX3080 graphics card. Future versions will be reimplemented using C++ with multi-threading to meet real-time performance and practical application requirements.

### v2

@@ IV-C Triangle Descriptors
Inspired by the STD (Stable Triangle Descriptor) [30], we leverage both corners and wall segments to construct triangle descriptors. These descriptors are used to establish correspondences between the BIM and the LiDAR submap.
Corners extracted from the LiDAR submap and the BIM are used to form a neighborhood graph, where vertices represent corners and edges connect two corners if they are within a maximum distance of $L_{max}$. Cliques in this graph with three vertices are used to form triangle descriptors, denoted by $\Delta$. These triangles are formed by corner triplets $(A,B,C)$, as illustrated in Figure 3. The attributes of a triangle descriptor $\Delta$ include:
@@
Figure 4: The figure demonstrates the effectiveness of triangle descriptors in matching results, with and without angle inclusion. On the left, a highlighted corner triplet from the LiDAR submap is shown in orange, which has multiple matches from the BIM using descriptor-based matching. Incorporating angle information significantly enhances the discriminative power of the triangle descriptor by encoding the local geometric structure around the triangle, which markedly reduces the false positives.
These sorted side lengths and angles capture both the geometric and structural characteristics of the local environment, distinguishing this approach from the original STD descriptor [30]. Furthermore, they are invariant to rotation and translation, ensuring independence from the coordinate systems of the BIM and submaps. As shown in Figure 4, querying each triangle in the submap can yield multiple responses in the BIM due to repetitive architectural patterns. The inclusion of angles $\alpha,\beta,\gamma$ helps to differentiate triangles that have identical shapes but differ in their angular relationships with adjacent wall segments, thus significantly reducing incorrect correspondences.
For efficient data management and retrieval, we construct a hash table for both LiDAR submap and the BIM. The keys are formed using the triangle descriptor $\Delta=[\|AB\|,\|BC\|,\|AC\|,\alpha,\beta,\gamma]$, with $\Delta\in\mathbf{R}^{6}$, and are quantized based on side length resolution $r_{\mathrm{s}}$ and angle resolution $r_{\mathrm{a}}$ as detailed in Table II and analyzed in Section VI-C1. Multiple corner triplets that yield the same hash key are stored together:
$\Delta\mathcal{S}$ =$=\{\Delta_{i}^{s}\mid\mathcal{F}(\Delta_{i}^{s})=\{(A,B,C)_{j}\}_{j=1}^{K_{i}^{s}}\}_{i=1}^{N^{s}},$
$\Delta\mathcal{B}$ =$=\{\Delta_{i}^{b}\mid\mathcal{F}(\Delta_{i}^{b})=\{(A,B,C)_{j}\}_{j=1}^{K_{i}^{b}}\}_{i=1}^{N^{b}},$
in which $\mathcal{F}(\Delta)$ retrieves the set of corner triplets $(A,B,C)$ corresponding to the descriptor key $\Delta$. The indices $N^{s}$ and $N^{b}$ denote the number of unique triangle descriptors in the submap and the BIM, respectively, while $K^{s}_{i}$ and $K^{b}_{i}$ represent the number of corner triplets associated with each descriptor $\Delta_{i}^{s}$ and $\Delta_{i}^{b}$.
All triangle descriptor keys and their associated corner triplets from the BIM are precomputed and stored in a hash database. During the registration stage, this database is loaded, and descriptors from the submaps are matched against the descriptors in the BIM. The matching process is performed using hash retrieval, which operates with a constant-time query complexity of $O(1)$. For each matched descriptor pair, all corresponding corner triplet candidates are exhaustively evaluated. Although this process generates an extremely large number of candidate correspondences with a high outlier ratio, it also retains a high likelihood of including the true correspondences. The next section will describe the method used to identify these true correspondences.
@@ V Back-end: Transformation Voting and Verification
In this section, Section V-A outlines the methodology for deriving pose transformation candidates from correspondences. This is achieved by using the Hough transform combined with a hierarchical voting mechanism. Subsequently, Section V-B introduces a verification scheme that identifies the optimal transformation candidate by employing an occupancy-aware confidence score to evaluate alignment quality, thereby ensuring reliable registration.
V-A Pose Hough Transformation and Hierarchical Voting
@@
The Hough transform is a robust technique in computer vision for converting raw measurements into a parameter space, where optimal parameters are identified through a hierarchical voting mechanism to effectively reject outliers.
V-A1 Parameter Space
The parameter space is based on the $\operatorname{SE}(2)$ Lie group, which represents planar rigid body transformations with 2 degrees of freedom (DOF) for translation and one for rotation (yaw). This space is discretized into voxels at resolutions defined by $r_{\mathrm{xy}}$ for x-y translation and $r_{\mathrm{yaw}}$ for yaw, as detailed in Table II and further discussed in the ablation study (Section VI-C2). For each corner triplet correspondence $\left[(A,B,C)_{s},(A,B,C)_{b}\right]$, transformations $(x,y,\text{yaw})$ are computed using a closed-form solution for 2D registration [31]. Each transformation contributes a vote for the corresponding voxel in the parameter space.
V-A2 Hierarchical Voting
Figure 5: The proposed hierarchical voting method consists of three steps: Broad Filtering, Neighborhood Refinement, and NMS with Ranking. Each step gradually reduces the number of transformation candidates, balancing computational efficiency and robustness.
To identify the most probable transformation candidates, we propose a hierarchical voting method that balances computational efficiency with robustness. The method comprises three steps:
  1. 1.
    Broad Filtering: The top $L$ voxels are selected through partial sorting based on their individual vote counts. This computationally efficient step reduces the dataset size while preserving a pool of potential transformation candidates for further processing.
  2. 2.
    Neighborhood Refinement: The $L$ selected voxels are refined to $K$ by merging votes within a $3\times 3\times 3$ neighborhood around each voxel. This step leverages local information to mitigate inaccuracies in corner triplet estimates, as votes may otherwise fall into adjacent voxels.
  3. 3.
    NMS and Ranking: A region-growing clustering algorithm is applied to the $K$ refined voxels, resulting in $K^{\prime}$ clusters. Non-maximum suppression (NMS) is achieved by ranking each cluster according to the maximum vote count of its constituent voxels. For each cluster, the transformation parameters are averaged to derive a representative transformation. Finally, the top $J$ clusters are selected as the most likely transformation candidates.
This hierarchical approach progressively reduces the candidate pool ($L>K>K^{\prime}>J$), achieving a balance between computational efficiency and recall. By incorporating neighborhood information, the method improves robustness and ensures reliable alignment results while maintaining manageable computational costs.
(a) More votes with lower confidence score
(b) Less votes with higher confidence score
Figure 6: Transformation verification. In each sub-figure, the first column displays the LiDAR submap and BIM along with their extracted corner triplets. The second column presents the score field $M_{d}$ as well as the ground ($Q_{g}$) and non-ground ($Q_{ng}$) components of the transformed LiDAR submap, are presented. In the third column, the alignment using the transformation candidate is shown, along with the corresponding confidence score. Figures 6(a) and 6(b) showcase two transformation candidates, where the confidence score successfully identifies the correct transformation candidate, whereas the votes do not. While both candidates align the walls and corners accurately, our proposed occupancy-aware confidence score penalizes the false candidate due to regions where the submap is free but the BIM is occupied.
@@ V-B Transformation Verification
Given $J$ transformation candidates, our objective is to develop a verification scheme to determine the optimal transformation which align the submap and BIM best. While corners construct the correspondences between the submap and BIM in previous sections, they don’t capture enough geometric structure for alignment quality evaluation. As illustrated in Figure 6, both transformation candidates can align the walls and corners, with the false candidate even receiving more votes from the matched corner triplets. To address this issue, we utilize the original submap point cloud to compute an occupancy-aware score. This score incorporates original and richer geometric information, offering a more robust evaluation of alignment quality to reliably identify the optimal transformation.
Denote the BIM wall point cloud as $P_{\text{B}}$, and the ground and non-ground planar components of the transformed LiDAR submap as $Q_{g}$ and $Q_{ng}$, respectively. An occupancy-aware confidence score is proposed to evaluate a transformation candidate by assessing the alignment quality between the BIM and the transformed submap using the given transformation.
Specifically, a rasterization function $\phi(\cdot,s_{r})$ is used to map $P_{\text{B}}$ to a 2D grid $M$ with a resolution of $s_{r}$, where occupied cells are set to 1 and free cells to 0:
$M=\phi(P_{\text{B}},s_{r}).$ (3)
To account for as-built deviations, the grid $M$ is dilated using a kernel of size $k_{d}$, producing the score field $M_{d}$. The value of each cell in $M_{d}$ decreases linearly with the distance to the nearest originally occupied cell, ranging from 1 at the nearest cell to $\frac{1}{k_{d}}$ at the kernel’s edge. This dilation introduces gradual transitions in cell values, enhancing robustness to noise and inaccuracies.
Award Component
The method rewards points for regions where the BIM wall point cloud and the actual construction $Q_{ng}$ match. The award score $s_{a}$ is calculated as:
$s_{a}=\sum_{\mathbf{q}\in Q_{ng}}M_{d}(\phi(\mathbf{q},s_{r})),$ (4)
where $M_{d}(\phi(\mathbf{q},s_{r}))$ denotes the value of the cell corresponding to the rasterized point $\phi(\mathbf{q},s_{r})$. We do not consider the situation where both the submap and the BIM are free, as this is less important than the occupied cases and may bias the score if included.
Penalty Component
Structures included in the BIM model are expected to be reflected in the actual construction. To penalize areas where the BIM is occupied but the submap is free, the penalty score $s_{p}$ is computed as:
$s_{p}=\sum_{\mathbf{q}\in Q_{g}}M_{d}(\phi(\mathbf{q},s_{r})).$ (5)
We do not consider the situation where the submap is occupied but the BIM is free, as real construction environments typically include additional elements such as lockers and furniture, which is a normal occurrence, especially when the BIM is outdated.
Final Confidence Score
The final confidence score is defined as:
$\text{score}=\frac{s_{a}-\lambda\cdot s_{p}}{|Q_{ng}|},$ (6)
where $\lambda$ is a hyperparameter that balances the contributions of the award and penalty terms. The score is normalized by dividing by $|Q_{ng}|$, the number of non-ground points in the transformed LiDAR submap, ensuring robustness to submaps of varying sizes. In the ideal case, when all points in $Q_{ng}$ match the BIM and there is no penalty from ground points, the score reaches its maximum value of 1.
This confidence score provides a quantitative and normalized measure of how effectively the transformation aligns the submap with the BIM. Furthermore, it can be utilized by users to assess the reliability of the registration, as demonstrated in Experiment VI-D.
@@ VI Experiments
This section is structured as follows: In Section VI-A, we introduce the dataset, evaluation metrics, and implementation details of our proposed method, as well as those of the methods against which we compare. Section VI-B presents comparative experiments that assess the registration recall of our method relative to other advanced methods. In Section VI-C, we explore how different parameters affect the performance of our proposed algorithm through an ablation study. In Section VI-D, we demonstrate the ability of our proposed occupancy-aware score to assess the reliability of the registration. Section VI-E provides visualizations of our algorithm applied in real-world scenarios and discusses cases where the algorithm failed.
@@ VI-A1 LiBIM-UST Dataset
TABLE I: LiBIM-UST Dataset Information
LiDAR Type Sequence Total Distance (m) Submaps (15 m) Submaps (30 m)
Ouster-128 Building day 402.14 169 -
Mid-360 1f-office 318.24 118 111
2f-office 533.78 193 171
3f-office 187.55 44 36
4f-office 191.52 76 69
5f-office 123.63 27 19
We developed the LiBIM-UST dataset using the publicly available FusionPortable dataset [32] and self-collected data for evaluation. The details are provided in Table I, where $d_{s}$, described in Section IV-A1, is set to different values for submaps of varying sizes. The sensor configuration for the self-collected data is illustrated in Figure 7. This dataset includes the BIM of the main academic building at the Hong Kong University of Science and Technology (HKUST). The BIM model is part of the digital twin project at HKUST, which was initiated in 2019.
The model was developed using Autodesk Revit and is stored in the .RVT file format with a Level of Detail (LoD) of 300 [33]. It was created based on archival 2D drawings from 1988, where the interior was primarily modeled from floor plans, and the exterior appearance was reconstructed using site photos and point cloud data. However, the model lacks textural information and contains several discrepancies when compared to the actual built environment as of 2023.
@@ VI-A3 Comparisons and Implementation
Image Registration: For image registration, intermediate images from the line extraction process (Section IV-B) are used. We implemented the SIFT [34] and a combination of Superpoint [35] with LightGlue [36], representing traditional and deep learning-based methods, respectively. RANSAC, with up to $1\times 10^{8}$ iterations, was used for pose estimation.
Point Cloud Registration: We employed three common global point cloud registration methods: FPFH [37] combined with TEASER++ [38], GeoTransformer [22], an advanced deep learning based registration method, and 3D-BBS [39], a branch-and-bound based method. For FPFH, we set the normal search radius to 1 m and the feature search radius to 2.5 m. For TEASER++, the noise bound is set at 0.4 m. For GeoTransformer, we retrain it on our datasets. For 3D-BBS, the parameter $\mathrm{min\_level\_res}$ is set to 0.2 m and the $\mathrm{max\_level}$ is set to 6.
Our Method: Detailed parameters of our method are listed in Table II. The BIM preprocessing and the Hash table configuration were predetermined. The experiments were conducted on an Intel i7-13700K CPU and an Nvidia RTX4080 graphics card.
TABLE II: Default Parameters of The Proposed Method
Parameters Value
Wall Extraction $r_{v}$ 0.8 m
$\sigma_{\lambda}$ 10
$s_{\mathrm{I}}$ 60 pixels / m
$L_{\mathrm{min}}$ 30 pixels
Triangle Descriptor $r_{\mathrm{s}}$ 0.5 m
$r_{\mathrm{a}}$ 3.0
$L_{\mathrm{max}}$ 30.0 m / 40.0 m
Hough Transform and Voting $r_{\mathrm{xy}}$ 0.15 m
$r_{\mathrm{yaw}}$ 1.0
L / K / J 10000 / 5000 / 1500
TABLE III: Global Registration Evaluation on LiBIM-UST Dataset
Front-end Back-end Recall (%)
BuildingDay 1F 2F 3F 4F 5F
$d_{s}$=15m $d_{s}$=15m $d_{s}$=30m $d_{s}$=15m $d_{s}$=30m $d_{s}$=15m $d_{s}$=30m $d_{s}$=15m $d_{s}$=30m $d_{s}$=15m $d_{s}$=30m
Image Matching- based Method SIFT RANSAC 0 0 0 0 0 0 0 0 0 0 0
Superpoint+Lightglue RANSAC 0 0 0 0 0 0 0 0 0 0 0
Point Cloud- based Method FPFH TEASER++ 0 0 0 1.04 0 0 0 1.32 0 0 0
GeoTransformer Weighted SVD 0 6.67 0 0 0 7.69 0 0 0 0 0
Voxel Map 3DBBS 86.39 41.52 47.74 75.65 94.74 100.0 100.0 67.10 94.20 100.0 100.0
Ours Triangle Descriptor Hough Voting 98.81 64.40 89.18 77.20 98.83 97.72 100.0 88.15 100.0 100.0 100.0
Front-end Back-end Average Time (ms)
BuildingDay 1F 2F 3F 4F 5F
$d_{s}$=15m $d_{s}$=15m $d_{s}$=30m $d_{s}$=15m $d_{s}$=30m $d_{s}$=15m $d_{s}$=30m $d_{s}$=15m $d_{s}$=30m $d_{s}$=15m $d_{s}$=30m
Image Matching- based Method SIFT RANSAC 6786.33 10670.65 8946.49 12935.78 13034.09 7083.24 5618.32 12683.18 14477.53 8283.46 5092.87
Superpoint+Lightglue RANSAC 499.93 859.68 1209.79 500.43 794.47 302.09 392.47 650.38 791.70 317.18 486.24
Point Cloud- based Method FPFH TEASER++ 25694.49 4169.82 4799.47 473.214 921.16 455.51 818.19 517.70 666.60 601.38 717.67
GeoTransformer Weighted SVD 441.3 492.7 275.0 378.9 397.3 427.5 366.7 415.7 432.2 467.1 422.9
Voxel Map 3DBBS 6603.00 2850.90 5284.33 2083.00 2862.00 821.72 1408.35 3658.24 4158.97 1373.10 1674.25
Ours Triangle Descriptor Hough Voting 121.34 36.72 95.14 54.96 170.12 77.42 369.60 43.30 152.83 68.79 268.59